OpenStudy (anonymous):

5y^3-2y^2+5y-2 factor

OpenStudy (anonymous):

(5y-2)(y^2+1)

OpenStudy (nathan917):

If you were to simplify it would be: Simplifying x + 3 + x + 4 = x + 7 + 2x + 7 Reorder the terms: 3 + 4 + x + x = x + 7 + 2x + 7 Combine like terms: 3 + 4 = 7 7 + x + x = x + 7 + 2x + 7 Combine like terms: x + x = 2x 7 + 2x = x + 7 + 2x + 7 Reorder the terms: 7 + 2x = 7 + 7 + x + 2x Combine like terms: 7 + 7 = 14 7 + 2x = 14 + x + 2x Combine like terms: x + 2x = 3x 7 + 2x = 14 + 3x Solving 7 + 2x = 14 + 3x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3x' to each side of the equation. 7 + 2x + -3x = 14 + 3x + -3x Combine like terms: 2x + -3x = -1x 7 + -1x = 14 + 3x + -3x Combine like terms: 3x + -3x = 0 7 + -1x = 14 + 0 7 + -1x = 14 Add '-7' to each side of the equation. 7 + -7 + -1x = 14 + -7 Combine like terms: 7 + -7 = 0 0 + -1x = 14 + -7 -1x = 14 + -7 Combine like terms: 14 + -7 = 7 -1x = 7 Divide each side by '-1'. x = -7 Simplifying x = -7